the lifetimes of a brand of electric light bulbs are normally distributed with mean 1300hours and standartd deviation 120hours. a light bulb is selected at random and tested by keeping it on continuously
find the probability that it has a lifetime of over 1540hours?
its has a lifetime between 1222 and 1540 hours?
two light bulbs are selected at random.
find the probability that they both last between 1222 and 1540

Answer by Boreal(15235)

z=(x-mean)/sd
=(1540-300/120=2
probability z>2=0.0228
-----------------
(1222-1300)/120=-78/120=-0.65
probability -0.65 -----------------
That would be the probability just obtained squared, since they are independent: 0.5175 if I don't round the last one.